A company that operates for 18 hours a day has introduced a new daily wage system that
incorporates variable hourly rate to encourage longer shifts. The company pays $10 for first
hour or any part of the first hour. Afterwards, company pays additional 2% per hour for each
hour or part of each additional hour. Write a piece-wise function that shows the salary as a
function of hours. Also sketch a graph of this function.
1
Expert's answer
2021-09-27T14:15:50-0400

Let "t" represent the hours and let "f(t)" be the salary as function of hours "t".

For the first hour or any part of the first hour;

"f(t)=10~~~0<t\leq1"

For the additional hours. That is "1<t\leq18"

"f(t)=10+(2\%\times10)(t-1)\\
f(t)=10+0.2t-0.2\\
f(t)=9.8+0.2t"

Hence the piece-wise function representing the salary is\\

"f(t)=\begin{cases} 10 ~~~~~~~~~~~~~~~~~~~~0<t\leq 1\\
9.8+0.2t ~~~~~~~~1<t\leq18 \end{cases}"

The graph is as follows